# Learning policies for resource allocation in business processes

### v2

@@ 3.2 Business process execution
The last part of the optimization problem is the objective, which is to minimize the average cycle time of cases in the process.
Definition 9 (Completed Case, Cycle Time)
@@
For a sequence of events $\sigma=\langle e_{1},e_{2},\dots,e_{n}\rangle$ of a case, the cycle time $c_{CT}$ is the difference between the time of occurrence of the first and the last event, i.e., $c_{CT}=\#_{t}(e_{n})-\#_{t}(e_{1})$. We can determine the cycle time for an ongoing case #$\#_{c}(e_{i})\in C$ or a completed case #$\#_{c}(e_{i})\in\mathcal{C}-C$.
Thus, the objective function can be written as:
$\operatorname*{minimize}\frac{1}{|\mathcal{C}-C|}\sum_{c\in\mathcal{C}-C}c_{CT}$ (1)
The interaction between the environment and the decision-making agent is as follows. In any given business process execution state (Def. 5), the environment checks for any possible assignments (Def. 7). If there are, the agent is asked to make a single assignment, and the environment transitions to the next business process execution state (Def. 5) by removing the respective available resource (Def. 4) and unassigned activity instance (Def. 3). If there are still possible assignments (i.e., $D\neq\emptyset$), the agent is asked for another assignment. When there are no more possible assignments or if the agent decides to postpone its decision, the business process simulation resumes, which modifies the business process execution state according to Def. 8. After the business process execution state changes, the environment checks for possible assignments again.
@@ 4 Method
In this section, we present two learning-based methods for resource allocation in business processes to minimize the average cycle time of cases. First, we present a Deep Reinforcement Learning (DRL) method, and second, we present a Score-based Value Function Approximation (SVFA) method.
The main differences between the two methods are how each algorithm is trained and how information is represented in each algorithm. The DRL method learns continuously during the simulation of the business process based on a reward it receives after each decision step, whereas SVFA receives a single reward signal at the end of the simulation. SVFA’s advantage is that it directly minimizes the objective function. Conversely, rewarding the agent to minimize the objective function in our DRL approach can introduce bias. For example, our reward function might prioritize the quick completion of one case while delaying others, leading to high rewards but detrimental effects on the average cycle time. The information provided in the SVFA method is represented in a set of hand-crafted features based on domain knowledge and optimizes these weights. On the other hand, the DRL agent uses a neural network that automatically learns important features from the state space. The hand-crafted features contain important information for resource allocation but may overlook vital information hidden in the data, which a neural network can discover.
We provide two methods to demonstrate different ways to approach resource allocation for business processes using learning-based methods. Section 4.1 details the DRL method and Section 4.2 the SVFA method.
@@ 4.1 Deep Reinforcement Learning
According to Sutton & Barto sutton_reinforcement_2018 , RL is a method to train one or multiple agents to interact with an environment to maximize a reward signal. The term DRL is used for RL algorithms that employ neural networks as function approximations. Fig. 2 shows how we developed the reinforcement cycle for resource allocation in business processes. The environment is a discrete-event simulation of the business process. Each time an assignment is possible, according to Def. 7, the simulation prompts the agent to make a decision, which is an assignment. We refer to this moment as a decision step. The environment processes the agent’s chosen action, and based on the action’s effect on the environment, a reward is returned to the agent. One run of the simulation is called an episode and can contain a different number of decision steps, depending on the stochastic events that take place in each run. The agent learns a policy, which is a function that maps the state to the actions and defines which action should be taken in a given state to maximize the reward.
Figure 2: Agent’s perspective of the agent-environment interaction, adapted from sutton_reinforcement_2018 .
In this paper, we use the state-of-the-art DRL algorithm Proximal Policy Optimization (PPO) schulman_proximal_2017 . Depending on the available resources and unassigned activity instances, several assignments may not be possible at a specific decision step (Def. 7). Therefore, we use a masking function that, given a state, prevents the agent from taking these infeasible actions, improving training speed and scalability huang_closer_2022 . We model the resource allocation problem as a finite Markov Decision Process (MDP) within the DRL framework. The remainder of this section defines each component of the MDP, as shown in Fig 2.
@@
The business process execution state (Def. 5) contains all information related to the current condition of the process. The state, as defined in the RL paradigm, describes what the agent observes from the business process execution state. To create the features in the state, we order the resources $r\in\mathcal{R}$ and activities $a\in\mathbf{A}$. We use subscript $r_{i}$ and $a_{j}$ to indicate the position of a specific resource and activity in the list, where $i\in\{1,\dots,|\mathcal{R}|\}$ and $j\in\{1,\dots,|\mathcal{A}|\}$. The state space contains:
    A mapping $\delta_{i}\rightarrow\{0,1\}$, indicating for each resource whether it is assigned or not (Def. 4), e.g., for resource $r_{i}:1$ if $r_{i}\in R^{+}$, or 00 if $r_{i}\in R^{-}$.
    A mapping $\eta_{i}\rightarrow[0,1]$, indicating for each resource which activity the resource is assigned to (Def. 7), e.g., for resource $r_{i}:\{\frac{j}{|\mathcal{A}|}\mid k\in K_{a_{j}}\}$, if $(r_{i},k)\in B$, or 0 otherwise.
@@
The input of the agent is the state vector. As shown above, the features in the vector are normalized between 0 and 1, which is a common practice to improve sample efficiency during training henderson_deep_2018 . Given a state, the agent receives information about which resources are available, how many unassigned activity instances are queued at each activity, and the current assignments. Using this information, the agent can learn what assignments lead to high rewards based on the state. Furthermore, it can learn when other resources are likely to be available again, and it can, for example, postpone a decision to wait for a better assignment. The size of the state vector is $2\cdot|\mathcal{R}|+|\mathcal{A}|$.
@@
The actions are all assignments (Def. 7) and a postpone action, which delays the current decision step and evolves the environment to the next decision step. At each decision step, a single action is taken. As mentioned, we mask infeasible actions ($\mathcal{D}-D$ according to Def. 7). The masking function is a vector that maps $\mathcal{D}\rightarrow\{0,1\}$, indicating if an assignment is possible or not, i.e., $1$ if $(r,k)\in D$ and 00 otherwise. The postpone action is always feasible. The action vector is multiplied by the masking vector, setting the probability of choosing infeasible actions to zero. The size of the action vector is $|\mathcal{D}|+1$.
@@
Given a specific state and action, the business process transitions to a new state, according to Def. 8. Multiple actions can be taken at the same moment in time, depending on the availability of resources or unassigned activity instances. In this case, actions are taken one at a time until no assignments are possible or the agent chooses the postpone action. If there are no possible assignments, the environment evolves until the next decision step. If the agent postpones, the environment evolves until there is at least one different unassigned activity instance or available resource compared to the previous state.
@@
The optimization objective is to minimize the average cycle time of cases (Def. 9), and we should reward the agent for actions that contribute to this objective. Therefore, we give a reward of $\frac{1}{c_{CT}+1}$ for each case $c$ that is completed. We inverse the cycle time such that a low cycle time results in a high reward and add $+1$ to the denominator to reduce the dimensionality of the rewards.
@@ 4.2 Score-based value function approximation
The second method we propose is a score-based value function approximation (SVFA) approach that, based on a set of features, learns to minimize the average cycle time of cases. We consider all possible assignments $(r,k)\in D$ (Def. 7) in a given business process execution state and compute a score for each of them. The score is a function of a set of features related to assignment $(r,k)$ and learned weights for each feature. The features are hand-crafted based on expert knowledge. The features are:
@@
To calculate the score, we aggregate the first four features in Equation 2 through addition because a lower value is better (i.e., lower cycle time). A high value for the last two features $ProbFin(r,k)$ and $QueueLength(k)$ is preferred, as finishing cases and minimizing queue lengths result in a lower average cycle time hopp2011factory . Therefore, these features are subtracted when calculating the score.
At each decision moment, the score for all possible assignments (Def. 7 is calculated. The assignment with the lowest score is the best and will be made unless its score does not exceed a learned threshold $w_{7}$. Instead, the agent postpones if no assignment’s score is lower than this seventh weight. Similarly to the DRL approach, when multiple assignments are possible, the agent makes assignments one at a time and then reevaluates all scores based on the new state to make the next assignment.
We use Bayesian optimization to find weights that minimize the average cycle time. Bayesian optimization is a global optimization technique that uses a probabilistic model to predict the relationship between the input parameters and the output of a function. The algorithm learns as new data is obtained and focuses on selecting parameters in promising areas of the search space. For more information on the Bayesian optimization algorithm, we refer to Frazier frazier2018tutorial . The implementation details are given in Section 5.2.2.
@@ 5.1 Experimental setup
This section presents the business processes on which we evaluate and benchmark our methods. First, we design six different business processes, each consisting of two activities and two resources, to analyze which methods work well in what specific situations. We refer to these six business processes as scenarios. Each scenario is characterized by process patterns such as parallel or choice constructs, varying utilization rates, and resource competition. Second, we connect these six scenarios sequentially and in parallel to create three realistically sized business processes containing all the different process flows seen in the scenarios. Evaluation of these business processes demonstrates which methods perform well under specific circumstances (i.e., the scenarios) and which methods perform well when all scenario dynamics are combined. We refer to these three business processes as composite business processes.
We first introduce the six scenarios. In each scenario, cases arrive according to a Poisson process with rate $\lambda=\frac{1}{2}$ time units. The processing times are exponentially distributed with mean $X_{r}$, $r\in\mathcal{R}$. The mean processing times are chosen to align with the characteristics of their respective scenario (e.g., utilization level, a slower resource) and ensure that each scenario’s cycle time is in the same order of magnitude. The latter is important for connecting the scenarios to create the composite business processes such that a single scenario does not dominate the total cycle time of the composite process. The six scenarios are shown in Fig. 3.
@@
Fig. 2(a) and Fig. 2(b) show the first two scenarios characterized by their utilization rates. The process’s utilization rate significantly impacts the average queue length at an activity hopp2011factory , i.e., $|K_{a}|$. If the utilization rate is low, such as in scenario 1 (Fig. 2(a)), the average size of the queues in the process will also be smaller, and therefore, the agent should focus on minimizing the processing time. Scenario 2 in Fig. 2(b) has a high utilization rate, which means that the average queue length will be higher, and cases will spend most of their cycle time queuing. Therefore, minimizing queueing time becomes more important in this scenario.
@@
Fig. 2(c) shows scenario 3 in which an experienced employee ($r_{5}$) works faster than an inexperienced employee ($r_{6}$). An efficient assignment policy should avoid assigning the inexperienced to activity $F$, as this resource is especially slow at executing it. Scenario 4 in Fig. 2(d) shows a business process where both resources have the same processing time distribution for each activity, of which the second activity, $H$, is processed slower than the first. Therefore, the prioritization of activities is more important than the resource that executes them.
@@
Fig. 2(e) models scenario 5, a process with two activities with different resource eligibility. This process is a variant of the well-studied bipartite matching problem (e.g., Cadas et al. cadas_bipartite_2019 ). It is also known as the ‘N-network’, referring to the ‘N’ shape of the bipartite graph, where nodes represent resources and activities, and edges represent the resource eligibility (Def. 6). The decision XOR gateway characterizes the process, where cases go to either activity with probability $p=0.5$. This scenario represents a situation where one employee ($r_{10}$) can perform more activities due to, for example, the authorization level. This resource’s capacity should be balanced over the activities. Fig. 2(f) shows scenario 6, which contains a parallel process flow. In this scenario, both activities can be performed simultaneously, but a case is only completed once both activities have been executed. The processing time at each activity is the same for both resources, making prioritizing what activity to complete the most important choice.
@@
In the second part of the evaluation, we create three realistically sized business processes by combining the six scenarios. We refer to these as composite business processes. The first composite business process is created by sequentially connecting the scenarios, starting with scenario 1 and ending with scenario 6. In the second composite business process, we also connect the scenarios sequentially but in reverse (i.e., starting with scenario 6 and ending with scenario 1) to investigate the impact of the order of the scenarios. Note that the external arrival process is still a Poisson process, while the internal arrivals now depend on departures from previous activities. The third composite business process is created by putting all scenarios in parallel.

### v3

@@ 3.3 Resource allocation problem
Note that the reward function depends on the optimization objective. In this paper, we aim to minimize the average cycle time of cases in the process.
Definition 14 (Cycle Time Reward)
@@
For a sequence of events of a case $\sigma=\langle e_{1},e_{2},\dots,e_{n}\rangle$, the cycle time $c_{CT}$ is the difference between the time of occurrence of the first and the last event, $c_{CT}=\#_{t}(e_{n})-\#_{t}(e_{1})$.
As the aim is to minimize the cycle time, a policy is optimal if it produces the lowest expected value.
Definition 15 (Optimal Resource Allocation Policy)
A policy $\pi$ is considered to be optimal if there is no other policy $\pi^{\prime}$ for which $\hat{V}_{\pi^{\prime}}(\Sigma)<\hat{V}_{\pi}(\Sigma)$ for a particular set of rollouts.
In practice, it is hard to determine if there exists no policy with a lower value. In addition to that, the policy that is selected to be optimal depends on the specific set of rollouts for which the values are determined. For that reason, the best policy is typically determined by approximation and subsequent comparison of the policy to other policies.
@@ 4 Method
In this section, we present two learning-based methods for resource allocation in business processes to approximate the optimal resource allocation policy that minimizes the average cycle time of cases (Def. 15). First, we present a Deep Reinforcement Learning (DRL) method, and second, we present a Score-based Value Function Approximation (SVFA) method.
The main differences between the two methods are how each algorithm is trained and how information is represented in each algorithm. The DRL method learns continuously during the simulation of the business process based on a reward it receives after each decision step, whereas SVFA receives a single reward signal at the end of the simulation. SVFA’s advantage is that it directly minimizes the objective function, while in DRL, an agent optimizes a feedback signal from a reward function. The challenge lies in developing a reward function that encourages the agent to optimize the objective.
The information provided in the SVFA method is represented in a set of hand-crafted features based on domain knowledge and optimizes these weights. On the other hand, the DRL agent uses a neural network that automatically learns important features from the state space. The hand-crafted features contain important information for resource allocation but may overlook vital information hidden in the data, which a neural network can discover.
Section 4.1 details the DRL method and Section 4.2 the SVFA method.
@@ 4.1 Deep Reinforcement Learning
According to Sutton & Barto [19], RL is a method to train one or multiple agents to interact with an environment to maximize a reward signal. The term DRL is used for RL algorithms that employ neural networks as function approximations. Figure 2 shows how we developed the reinforcement cycle for resource allocation in business processes. The environment is a discrete-event simulation of the business process. Each time an assignment is possible, according to Def. 8, the simulation prompts the agent to make a decision, which is an assignment. We refer to this moment as a decision step. The environment processes the agent’s chosen action, and based on the action’s effect on the environment, a reward is returned to the agent (Def. 13). One run of a simulation, $\Sigma$, is called an episode and can contain a different number of decision steps, depending on the stochastic events in each rollout. The agent learns a policy, which is a function that maps the state to the actions and defines which action should be taken in a given state to maximize the reward.
Figure 2: Agent’s perspective of the agent-environment interaction, adapted from [19].
In this paper, we use the state-of-the-art DRL algorithm Proximal Policy Optimization (PPO) [26]. Depending on the available resources and unassigned activity instances, several assignments may not be possible at a specific decision step (Def. 8). Therefore, we use a masking function that, given a state, prevents the agent from taking these infeasible actions, improving training speed and scalability [27].
The main difference between our DRL method and the most relevant previous work on DRL [8] is that we optimize for a more complex objective. Instead of only maximizing the number of completed cases, we also consider the cycle time in which they are completed. Furthermore, we mask infeasible actions, which significantly improve training speed. For example, in our largest process, there are 24 possible actions. If only one assignment is possible, [8] must take each action multiple times for every state to learn the feasible actions, while our method receives this information through the masking function. In the former method, there is still no guarantee infeasible actions cannot be taken. Lastly, we use PPO for our problem because it provides a stable and efficient learning process with gradual policy adjustments, which are crucial for handling the dynamic environment of online resource allocation.
We model the resource allocation problem as a finite Markov Decision Process (MDP) within the DRL framework. The remainder of this section defines each component of the MDP, as shown in Fig 2.
@@
The business process execution state (Def. 7) contains all information related to the current condition of the process. The state, as defined in the RL paradigm, describes what the agent observes from the business process execution state. To create the features in the state, we order the resources $r\in\mathcal{R}$ and activities $a\in\mathbf{A}$. We use subscript $r_{i}$ and $a_{j}$ to indicate the position of a specific resource and activity in the list, where $i\in\{1,\dots,|\mathcal{R}|\}$ and $j\in\{1,\dots,|\mathcal{A}|\}$. The state space contains:
    A mapping $\delta_{i}\rightarrow\{0,1\}$, indicating for each resource whether it is assigned or not (Def. 5), e.g., for resource $r_{i}:1$ if $r_{i}\in R^{+}$, or 00 if $r_{i}\in R^{-}$.
    A mapping $\eta_{i}\rightarrow[0,1]$, indicating for each resource which activity the resource is assigned to (Def. 8), e.g., for resource $r_{i}:\{\frac{j}{|\mathcal{A}|}\mid k\in K_{a_{j}}\}$, if $(r_{i},k)\in B$, or 0 otherwise.
@@
The input of the agent is the state vector. As shown above, the features in the vector are normalized between 0 and 1, which is a common practice to improve sample efficiency during training [28]. Given a state, the agent receives information about which resources are available, how many unassigned activity instances are queued at each activity, and the current assignments. Using this information, the agent can learn what assignments lead to high rewards based on the state. Furthermore, it can learn when other resources are likely to be available again, and it can, for example, postpone a decision to wait for a better assignment. The size of the state vector is $2\cdot|\mathcal{R}|+|\mathcal{A}|$.
@@
The actions are all assignments (Def. 8) and a postpone action, which delays the current decision step and evolves the environment to the next decision step. At each decision step, a single action is taken. As mentioned, we mask infeasible actions ($\mathcal{D}-D$ according to Def. 8). The masking function is a vector that maps $\mathcal{D}\rightarrow\{0,1\}$, indicating if an assignment is possible or not, i.e., $1$ if $(r,k)\in D$ and 00 otherwise. The postpone action is always feasible. The action vector is multiplied by the masking vector, setting the probability of choosing infeasible actions to zero. The size of the action vector is $|\mathcal{D}|+1$.
@@
Given a specific state and action, the business process transitions to a new state, according to Def. 9. Multiple actions can be taken at the same moment in time, depending on the availability of resources or unassigned activity instances. In this case, actions are taken one at a time until no assignments are possible or the agent chooses the postpone action. If there are no possible assignments, the environment evolves until the next decision step. If the agent postpones, the environment evolves until there is at least one different unassigned activity instance or available resource compared to the previous state.
@@
The optimization objective is to minimize the average cycle time of cases (Def. 14), and we should reward the agent for actions that contribute to this objective. As the agent maximizes cumulative discounted rewards, we inverse the cycle time reward (i.e., Def. 13) and give a reward of $\frac{1}{c_{CT}+1}$ for each case $c$ that is completed. We inverse the cycle time such that a low cycle time results in a high reward and add $+1$ to the denominator to ensure the rewards do not explode. The reward of a case completion is credited to the action that triggered the state transition in which the case completion happened. Other actions receive a reward of 0.
@@ 4.2 Score-based value function approximation
The second method we propose is a score-based value function approximation (SVFA) approach that, based on a set of features, learns to minimize the value of a policy (Def. 13). We consider all possible assignments $(r,k)\in D$ (Def. 8) in a given business process execution state and compute a score for each of them. The score is a function of a set of features related to assignment $(r,k)$ and learned weights for each feature. The features are hand-crafted based on expert knowledge. The features are:
@@
To calculate the score, we aggregate the first four features in Equation 1 by addition, as a lower value indicates better performance (i.e., lower cycle time). Conversely, higher values for the last two features, $ProbFin(r,k)$ and $QueueLength(k)$, are desirable because they signify efficient case completion and minimized queue lengths, which contribute to a lower average cycle time [29]. Therefore, these features are subtracted when calculating the score.
At each decision step, the score for all possible assignments (Def. 8) is calculated. The assignment with the lowest score is the best and will be made unless its score does not exceed a learned threshold $w_{7}$. Instead, the agent postpones if no assignment’s score is lower than this seventh weight. Similarly to the DRL approach, when multiple assignments are possible, the agent makes assignments one at a time and then reevaluates all scores based on the new state to make the next assignment.
We use Bayesian optimization to find weights that minimize the average cycle time. Bayesian optimization is a global optimization technique that uses a probabilistic model to predict the relationship between the input parameters and the output of a function. The algorithm learns as new data is obtained and focuses on selecting parameters in promising areas of the search space. In our setting, the input parameters are the feature weights, and the output is the mean cycle time after an episode. In contrast to our DRL approach, SVFA does not learn from the decision steps in an episode but updates the policy at the end of the episode, based on the average cycle time following the current policy. For more information on the Bayesian optimization algorithm, we refer to Frazier [30]. The implementation details are given in Section 5.2.2.
@@ 5.1 Experimental setup
This section presents the business processes on which we evaluate and benchmark our methods. First, we design six different business processes, each consisting of two activities and two resources, to analyze which methods work well in what specific situations. We refer to these six business processes as scenarios. Each scenario is an abstraction of pattern commonly seen in business processes. The scenarios are characterized by aspects such as parallel or choice constructs, varying utilization rates, and competition over resources. Second, we connect these six scenarios sequentially and in parallel to create three realistically sized business processes containing all the different process flows seen in the scenarios. Evaluation of these business processes demonstrates which methods perform well under specific circumstances (i.e., the scenarios) and which generalize well to a process in which all scenario dynamics are combined. We refer to these three business processes as composite business processes.
We first introduce the six scenarios. In each scenario, cases arrive according to a Poisson process with rate $\lambda=\frac{1}{2}$ time units. This arrival rate ensures the system remains stable while still offering a load on the resources such that the problem is challenging. In Section 5.4, we show that the processes become unstable when using higher arrival rates. The processing times are exponentially distributed with mean $X_{r}$, $r\in\mathcal{R}$. The mean processing times are chosen to align with the characteristics of their respective scenario (e.g., utilization level, a slower resource) and ensure that each scenario’s cycle time is in the same order of magnitude. The latter is important for connecting the scenarios to create the composite business processes such that a single scenario does not dominate the total cycle time of the composite process. The six scenarios are shown in Figure 3.
@@
Figure 3(a) and Figure 3(b) show the first two scenarios characterized by their utilization rates and represent situations in which the process operates under different loads, such as peak and off-peak hours. The process’s utilization rate significantly impacts the average queue length at an activity $|K_{a}|$ [29]. If the utilization rate is low, such as in scenario 1 (Figure 3(a)), the average size of the queues in the process will also be smaller, and therefore, the agent should focus on minimizing the processing time. Scenario 2 in Figure 3(b) has a high utilization rate, which means that the average queue length will be higher, and cases will spend most of their cycle time queuing. Therefore, minimizing queueing time becomes more important in this scenario.
@@
Figure 3(c) shows a variant of the slow server problem which is challenging and well-studied in the literature [31]. With respect to business process, this scenario contains an experienced employee ($r_{5}$) who works faster than an inexperienced employee ($r_{6}$). An efficient assignment policy should avoid assigning the inexperienced to activity $F$, as this resource is especially slow at executing it. Scenario 4 in Figure 3(d) shows a business process where both resources have the same processing time distribution for each activity. Therefore, the prioritization of activities is more important than the resource that executes them. This problem is relevant for processes with standardized processing times.
@@
Figure 3(e) models scenario 5, a process with two activities with different resource eligibility. This process is a variant of the well-studied bipartite matching problem [32]. It is also known as the ‘N-network’, referring to the ‘N’ shape of the bipartite graph, where nodes represent resources and activities, and edges represent the resource eligibility (Def. 6). The exclusive gateway (XOR) characterizes the process, where cases go to either activity with probability $p=0.5$. This scenario represents a situation where one employee ($r_{10}$) can perform more activities due to, for example, the authorization level. This resource’s capacity should be balanced over the activities. Figure 3(f) shows scenario 6, which contains a parallel process flow typically seen in business processes. In this scenario, both activities can be performed simultaneously, but a case is only completed once both activities have been executed. The processing time at each activity is the same for both resources, making prioritizing what activity to complete the most important choice.
@@
In the second part of the evaluation, we trained and tested our models on the three composite processes. The first composite business process is created by sequentially connecting the scenarios, starting with scenario 1 and ending with scenario 6. In the second composite business process, we also connect the scenarios sequentially but in reverse to investigate the impact of the order of the scenarios. Note that in these processes, the external arrival process is still a Poisson process, while the internal arrivals now depend on departures from previous activities. The third composite business process is created by putting all scenarios in parallel.
These scenarios provide typical resource-allocating problems that require adaptive methods to solve them. While these scenarios can be extended by, for example, incorporating additional resources or activities, they provide a good foundation for understanding how different algorithms perform in specific conditions, specifically different control flow patterns, resource loads, and authorization structures (such as the N-model). Moreover, applying our methods to the composite models allows us to study which methods can handle all scenario dynamics simultaneously.